iiXilosis:

The graph of a line goes through the points (-4,3) and (6,8). What is the equation of the line in slope-intercept form? Enter the correct answer in the box by replacing m and b with the appropriate values. y=mx+b

snowflake0531:

first find the slope Slope~Formula\[m=\frac{y_{2} - y_{1}}{x_{2}-x_{1}}\] in this case, \(m=\frac{8-3}{6-(-4)}=\frac{5}{10}=\frac{1}{2}\) now let's find b with the slope and one coordinate point m=1/2 (6,8) 8=1/2 (6)+b 8=3+b b=5 \[y=\frac{1}{2}x + 5\]